Circular No. 6040

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Postal Address: Central Bureau for Astronomical Telegrams Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A. Telephone 617-495-7244/7440/7444 (for emergency use only) TWX 710-320-6842 ASTROGRAM CAM EASYLINK 62794505 MARSDEN@CFA or GREEN@CFA (.SPAN, .BITNET or .HARVARD.EDU)

JUPITER AND COMET 1993e M. Kesteven and R. M. Price, Australia Telescope National Facility

(ATNF); I. de Pater, University of California at Berkeley; G. A. Dulk, University of Colorado; and Y. Leblanc, Observatory of Paris, write: “Observations of Jupiter were made in Australia at a frequency of 1430 MHz: at Parkes (64-m dish, July 15–24) and with the compact array of the ATNF (6 dishes of 22 m each, July 9–25). These observations show an increase in the flux density of ∼ 0.5 Jy from a pre-impact level of ∼ 3.7 Jy. The major part of the increase occurred during July 17–20; subsequently, the level has been nearly steady. No change has been seen in the position angle of the linear polarization. The images produced by the ATNF compact array suggest that the increase in flux was due mainly to an increase of brightness in the inner, synchrotron-emitting belts, and not in the thermal emission of the disk.”

W. M. Harris, University of Michigan at Ann Arbor (UM); T. A. Livengood, Goddard Space Flight Center (GSFC); M. A. McGrath, Space Telescope Science Institute; R. Prange, (IAS), Orsay; G. E. Ballester, UM; S. Budzien, Naval Research Laboratory; M. Combi, UM; C. Emerich, IAS; G. Fireman, GSFC; D. T. Hall, Johns Hopkins University; K. L. Jessup, UM; A. Talavera, European Space Agency, Vilspa; M. B. Vincent, UM; and L. Woodney, University of Maryland, communicate: “We have analyzed spectra (range 115–195 nm) taken with the International Ultraviolet Explorer (IUE) of the Io torus and two impact regions during the recent impacts with Jupiter. Our early analysis of Io torus spectra taken on July 20 indicates that the torus was not significantly affected by initial phases of the collisions. We also monitored the temporal evolution of the E and G impact sites 40 min, 2 hr, and 4 hr after the impacts occured, as they rotated with the planet. There are clear changes in the spectrum consistent with the addition of new absorbers, including possibly ammonia and some hydrocarbon species, along with an enhancement in the strength of many previously-detected absorption features. We also detect an average 50-percent drop in the ultraviolet reflectivity of these regions across the IUE spectral bandpass, as they move from the dawn to the dusk limb of the planet. This result is consistent with the rapid development of a cloud of high-altitude absorbing material with an area close to the size of the IUE small (3′′) aperture, which at Jupiter is roughly equal that of the earth.”

1994 July 29 Daniel W. E. Green